OpenStudy (anonymous):

Evaluate. fourth root of 9 multiplied by square root of 9 over the fourth root of 9 to the power of 5 5 to the power of negative 1 over 6 5 to the power of 3 over 2 9 92

OpenStudy (anonymous):

@studygurl14

OpenStudy (anonymous):

Im so lost

OpenStudy (studygurl14):

This? \(\huge\frac{\sqrt[4]{9}\sqrt{9}}{(\sqrt[4]{9})^5}\)

OpenStudy (anonymous):

Yes

OpenStudy (studygurl14):

\(\large\sqrt[4]{9}=9^{\frac{1}{4}}\) And... \(\large(\sqrt[4]{9})^5=9^{\frac{5}{4}}\)

OpenStudy (studygurl14):

does that help?

OpenStudy (anonymous):

im still really confused now would i subtract the powers?

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (studygurl14):

oh, and \(\large\sqrt{9}=9^{\frac{1}{2}}\)

OpenStudy (anonymous):

I think the answer is D

OpenStudy (studygurl14):

So you're left with... \(\huge\frac{9^{\frac{1}{4}}9^{\frac{1}{2}}}{9^{\frac{5}{4}}}\)

OpenStudy (anonymous):

nvm

OpenStudy (studygurl14):

simplfy that

OpenStudy (anonymous):

How?

OpenStudy (studygurl14):

Do this: \(\huge 9^{\frac{1}{4}+\frac{1}{2}-\frac{5}{4}}\)

OpenStudy (anonymous):

so the base will stay 9 and in the answers the base is always 9 it just copied it wrong. and it would be 9 -1/-2

OpenStudy (anonymous):

as @StudyGurl14 explained 9^(1/4+1/2-5/4)=9^-2/4=9-1/2

OpenStudy (studygurl14):

It would be \(\large 9^{-\frac{1}{2}}\)

OpenStudy (anonymous):

thats not an answer choice tho

OpenStudy (anonymous):

Oh yes it is thank you

OpenStudy (studygurl14):

:)

OpenStudy (studygurl14):

can i have a medal now @Helppme6

OpenStudy (anonymous):

\[\frac{ 1 }{ \sqrt{9} }=\frac{ 1 }{ 3}\]